ci &si
In How many years will a sum of money trebles @5% p.a compound interest payable on half yearly basis ans:22 years 3 months

Answer by addingup(3677)

Since interest is payable 2 times per year, all calculations are on that basis.
We have to use the formula for n (number of periods) which is derived from the traditional compound interest formula FV= PV(1+r/n)^n. When you work the formula, you have to use logs to get rid of the ^n and you get:
------------------------------------
n= (ln(3000/1000))/(ln(1 + .025))
n=
1.099/.0247= 44.494/2= 22.247 and .297*12= 2.964 round to 3 and you get 22 years and 3 months.
J